OpenStudy (zephyr141):

help with an exam study guide question! i've never missed a class and never seen this question or went into more detail about it. I didn't see this on any homework assignments but could someone give me the rundown on this question?

OpenStudy (zephyr141):

The Coulomb Force is the force between two electrical charges. The potential energy associated with this force is\[U_{coulomb}(0)=\frac{1}{4\pi \epsilon_{0}}\frac{q_{1}q_{2}}{x}\]and epsilon(0), q(1) and q(2) are constants. Find the Coulomb force. a)\[F_{coulomb}(x)=\frac{1}{4\pi \epsilon_{0}}q_{1}q_{2}\ln x \]b)\[F_{coulomb}(x)=\frac{1}{4\pi \epsilon_{0}}\frac{q_{1}q_{2}}{x^2}\]c)\[F_{coulomb}(x)=\frac{-1}{4\pi \epsilon_{0}}\frac{q_{1}q_{2}}{x^2}\]d)\[F_{coulomb}(x)=\frac{-2}{4\pi \epsilon_{0}}\frac{q_{1}q_{2}}{x^3}\]

OpenStudy (anonymous):

Energy is measured in units of Joules. A [ J ] is a [N m]. Force is measured in units of Newtons. So a force is just energy divided by a displacement in meters. If the potential energy function has 1/x dependence, then the force would have (1/x) / (x) = 1/(x^2) dependence.